OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -3 + 6 + 15 + 24 + ... + 132

OpenStudy (solomonzelman):

Lets do this. can you tell me the difference ?

OpenStudy (solomonzelman):

oops the common ratio

OpenStudy (solomonzelman):

the common ratio is ?

OpenStudy (anonymous):

=9?

OpenStudy (solomonzelman):

Yes, it is difference

OpenStudy (solomonzelman):

The sequence confused me, sorry. Yes d=9

OpenStudy (solomonzelman):

How many terms ?

OpenStudy (solomonzelman):

\(\large\color{blue}{ a_n=a_1+d(n-1)}\) solving for n \(\large\color{blue}{ 132=-3+9(n-1)}\) \(\large\color{blue}{ n=?}\)

OpenStudy (anonymous):

n=23?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

my bad

OpenStudy (solomonzelman):

add 3 to both sides divide both sides by 9 add 1 to both sides it is not 23

OpenStudy (solomonzelman):

15 is `n-1` and `n` = ?

OpenStudy (anonymous):

16 lol

OpenStudy (solomonzelman):

Yes

OpenStudy (solomonzelman):

So a1=3 d=9 number of terms =16

OpenStudy (solomonzelman):

Fill in the Σ (alt code for Σ, is 2 2 8 )

OpenStudy (solomonzelman):

\[\LARGE \sum_{index}^{number~~of~~terms}~~(equation)\]

OpenStudy (solomonzelman):

index is what you start putting in for "n" into the equation to get whatever term.

OpenStudy (solomonzelman):

Go, shoot !

OpenStudy (anonymous):

15 Σ-27n n=0

OpenStudy (anonymous):

i tried :/

OpenStudy (anonymous):

or is it -3+9n?

OpenStudy (anonymous):

instead of multiplying

OpenStudy (solomonzelman):

would \[\sum_{n=0}^{15}~~(-3+9n)\]suffice ?

OpenStudy (anonymous):

yes, that was my second guess

OpenStudy (anonymous):

i didn't know which one

OpenStudy (solomonzelman):

Yes:)

OpenStudy (solomonzelman):

I want to outline what we did

OpenStudy (anonymous):

go for it

OpenStudy (solomonzelman):

1) Found the difference 2) Found the number of term 3) Found a sum notation would you be able to do it on your own in the future ?

OpenStudy (solomonzelman):

2) ... *terms

OpenStudy (anonymous):

i think so

OpenStudy (solomonzelman):

Good luck with your next question.... !

OpenStudy (anonymous):

wait, is there something i did wrong?